OpenStudy (anonymous):

5a+2b=-28 -5a+b=46 What is the solution?

OpenStudy (anonymous):

What I would do is add 5a to the bottom equation which will turn out as b=5a+46. Then I will plug it in into the top which the equation will look like 5a+2(5a+46)=-28. And then do the 2(5a+46) which is 10a+92. Now the equation will look like this, 5a+10a+92=-28 and then subtract 92 on both sides.then divide by 15

OpenStudy (wolf1728):

I think there is an easier way to do this: A) 5a+2b=-28 B) -5a+b=46 Add the 2 equations which results in 3b = 18 b = 6 Now you already have one of the unknowns solved. "Plug" the value of b (which is 6) in to either Equation A or B to get the value of a.

OpenStudy (anonymous):

yup that works too